OpenStudy (anonymous):

4 3/5+8 1/2+7 2/3=? show steps please

OpenStudy (ghazi):

\[4\frac{ 3 }{ 5 }= 4*5+3/5= \frac{ 23 }{5 }\] similarly 8 1/2= 17/2 and 7 2/3=23/3 therefore \[\frac{ 23 }{ 5 }+\frac{ 17 }{ 2 }+\frac{ 23 }{ 3 }\]...now you can do this

OpenStudy (ghazi):

can you?